OpenStudy (anonymous):

Help !! Will fan and medal !! 5x-6y+52-x-12+2z A.6x-18y+7z B.4x-18y+7x C.4x-6y+7z D.6x-6y+7z

OpenStudy (kendricklamar2014):

combine like terms

OpenStudy (kendricklamar2014):

5x- x = ?

OpenStudy (anonymous):

5x

OpenStudy (kendricklamar2014):

nope, 5x - 1x = ?, x alone means 1x

OpenStudy (anonymous):

4x....?

OpenStudy (kendricklamar2014):

yes, now 52 - 12= ?

OpenStudy (anonymous):

40...?

OpenStudy (kendricklamar2014):

yes, now were done so u get B

OpenStudy (anonymous):

Ohhh hehe Thanks Luv !! You really helped me..how do you know how to do this??:))

OpenStudy (kendricklamar2014):

Studying and paying attention in class

OpenStudy (anonymous):

Nice !! I get it,I study too but its hard that im virtual school its very challenging hehe:))

OpenStudy (kendricklamar2014):

yah, and also taking notes

OpenStudy (anonymous):

Alot of notes Are you in Flvs?

OpenStudy (kendricklamar2014):

No

OpenStudy (anonymous):

Oh Sowvie hehe:)) thanks for helping me out :))

OpenStudy (kendricklamar2014):

No problem

OpenStudy (anonymous):

@KendrickLamar2014 If i happen to need more assistance would you be able to help me??:))

OpenStudy (kendricklamar2014):

yes

OpenStudy (anonymous):

Thank you soo much hehe:))..i'll just tag you whenever i need help :D

OpenStudy (kendricklamar2014):

ok